OpenStudy (anonymous):

((siny)+(tany))/((1)+(secy))=(siny)

OpenStudy (mathmale):

Kent: What is the objective of this problem? Could you possibly type the problem in using the Equation Editor? (If not, y ou've done better than average in presenting this problem clearly using parentheses.)

OpenStudy (anonymous):

how do i prove it i just need the left side to look the same as the right side

OpenStudy (mertsj):

\[\frac{\sin y+\tan y}{1+\sec y}=\sin y\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yeah

OpenStudy (jdoe0001):

\(\bf \cfrac{sin(y)+tan(y)}{1+sec(y)}=sin(y)\implies \cfrac{\frac{sin(y)}{1}+\frac{sin(y)}{cos(y)}}{\frac{1}{1}+\frac{1}{cos(y)}}\)

OpenStudy (mertsj):

1. Change tan y to siny/cosy 2. Change sec y to 1/cos y 3. Multiply numerator and denominator by cos y 4. factor sin y out of the numerator 5. cancel the common factor of numerator and denominator

OpenStudy (mathmale):

JDoe has shown nicely that numerator and denominator of the MAIN fractional expression have the same LCD. Can you, Kent, use that info to simplify that expression?